OpenStudy (anonymous):

Solve the following system of equations algebraically: 5x - 3y = 6 5x - 4y = 3

OpenStudy (shamil98):

Multiply the bottom equation by -1, and then add them.

OpenStudy (shamil98):

5x - 3y = 6 -5x + 4y = -3 what are you left with when you add this?

OpenStudy (anonymous):

Use elimination method y = 3 5x - 3(3) = 6 5x -9 = 6 5x = 15 x = 3 y = 3

OpenStudy (anonymous):

a.3,3 b.2,4 c.(3,3) and (0,-3) d. none are the answer choeses

OpenStudy (anonymous):

im mean choses

OpenStudy (anonymous):

It's A, I explain why in my previous comment.

OpenStudy (anonymous):

thank you